Massachusetts Institute of Technology Handout 2 6.042J/18.062J: Mathematics for Computer Science Feb 14, 1997 Professor David R. Karger

Tentative Rotation (off by 1 per week)

Problems are due by 5pm on Sunday. Problem set should be done by Monday morning, reviewed by Monday afternoon, and Prof Karger will do a final pass Monday night.

Problem set solutions should be done by Monday, reviewed on Tuesday, and given to Prof Karger on Wednesday.

{FEB \atop 3} 4

Lecture 1: – PSET 1 out.

5

6 Lecture 2:_{–}

7Recitation 1

FEB 10 Tutorial 1 11 Lecture 3: – PSET 1 due.

PSET 2 out.

12

13 Lecture 4:_{–}

14 Recitation 2

FEB 17

President’s Day. PSET 3–Alex, Daniele PSET 2 soln’s–Daniele, Alex 18 Monday Classes

Tutorial PS 2 Soln’s review–Dave PSET 2 due. PSET 3 out.

19 20 Lecture 5:_{Terri}

21 Recitation 3

FEB 24

Tutorial PSET 4– Terri, David PSET 3 soln’s–Alex, Daniele 25 Lecture 6:

Terri

PS 3 Soln’s review–Terri PSET 3 due. PSET 4 out.

26 27 Lecture 7:_{Alex} 28 Recitation 4 MAR 3

Tutorial PSET 5–Dave, Dimitri PSET 4 soln’s–Terri, David 4 Lecture 8:

Dave

PS 4 Soln’s review–Dimitri PSET 4 due. PSET 5 out.

5 6 Lecture 9:_{Daniele}

7Recitation 5

MAR 10

Tutorial PSET 6– Dimitri, Terri PSET 5 soln’s–Dave, Dimitri 11 Lecture 10:

Daniele

PS 5 Soln’s review– Danielle PSET 5 due. PSET 6 out.

12 13 Lecture 11:_{Dimitri}

14 Recitation 6

MAR 17

Tutorial PSET 6 soln’s–Dimitri, Terri 18 Lecture 12:

Dave

PS 6 Soln’s review–Alex PSET 6 due.

{19 \atop Quiz 1} 20

Lecture 13:_{Alex}

21 Recitation 7

APRIL 31

Tutorial PSET 7–Terri, Dimitri 1 Lecture 14: Dave PSET 7 out.

2 3 Lecture 15:_{Alex}

4Recitation 8

APRIL 7

Tutorial PSET 8–Dave, Dimitri PSET 7 soln’s–Terri, Dimitri 8 Lecture 16:

Dimitri

PS 7 Soln’s review–David PSET 7 due. PSET 8 out.

9 10 Lecture 17:_{Terri} 11 Recitation 9 APRIL 14

Tutorial PSET 8 soln’s–Dave, Dimitri 15 Lecture 18:

Terri(?)

PS 8 Soln’s review–Alex PSET 8 due.

16 17 Lecture 19:_{Daniele} 18

Recitation 10

APRIL 21

Patriot’s Day. PSET 9–Alex, Daniele 22

Patriot’s Day.

{23 \atop Quiz 2} 24

Lecture 20: Alex PSET 9 out.

25

Recitation 11

APRIL 28

Tutorial PSET 10– Daniele, Alex PSET 9 soln’s–Alex, Daniele 29 Lecture 21:

Daniele

PS 9 Soln’s review–Terri PSET 9 due. PSET 10 out.

30 1 Lecture 22:_{Alex} 2

Recitation 12

MAY 5

Tutorial PSET 11– David, Terri PSET 10 soln’s–Alex, Daniele 6 Lecture 23:

Daniele

PS 10 Soln’s review–Dimitri PSET 10 due. PSET 11 out.

7 8 Lecture 24:_{Dimitri} 9

Recitation 13

MAY 12

Tutorial PSET 11 soln’s–David, Terri 13 Lecture 25:

Terri

PS 11 Soln’s review– Daniele(?)

PSET 11 due.

14 15 Lecture 26:_{Dave(?)} 16